Given that A = 0.002 m². What pressure is exerted on the fluid F = 1000 N?
1
Expert's answer
2018-05-15T15:33:08-0400
Pressure is defined as force per unit area.

Pressure=Force/Area=F/A
Because pressure is force divided by area, its meter-kilogram-second (MKS) units are newtons per square meter, or N/m2. The unit newtons per square meter is so common in physics that it has a special name: the pascal, which equals 1 newton per square meter. The pascal is abbreviated as Pa.
So,
P=1000/0.002=500000 Pa